# The FLAMINGO project: cosmology with the redshift dependence of weak gravitational lensing peaks

### v1

@@ Abstract
Weak gravitational lensing (WL) convergence peaks contain valuable cosmological information in the regime of non-linear collapse. Using the FLAMINGO suite of cosmological hydrodynamical simulations, we study the physical origin and redshift distributions of the objects generating WL peaks selected from a WL convergence map mimicking a Euclid signal. We match peaks to individual haloes and show that the high signal-to-noise ratio (SNR > 5) WL peaks measured by Stage IV WL surveys primarily trace $M_{\mathrm{200c}}>10^{14}~{}\mathrm{M_{\odot}}$ haloes. We find that the WL peak sample can compete with the purity and completeness of state-of-the-art X-ray and Sunyaev-Zel’dovich cluster abundance inferences. By comparing the distributions predicted by simulation variations that have been calibrated to the observed gas fractions of local clusters and the present-day galaxy stellar mass function, or shifted versions of these, we illustrate that the shape of the redshift distribution of SNR > 5 peaks is insensitive to baryonic physics while it does change with cosmology. The difference highlights the potential of using WL peaks to constrain cosmology. As the number density of WL peaks scales differently with cosmology and baryonic feedback, WL peak statistics can simultaneously calibrate baryonic feedback and constrain cosmology.
@@ 1 Introduction
The cornerstone of modern cosmology is the $\Lambda$CDM model, consisting primarily of dark energy ($\Lambda$) and cold dark matter (CDM). The model can predict a wide range of cosmological observables, including the cosmic microwave background (CMB, e.g. Planck Collaboration et al., 2020a), baryonic acoustic oscillations (BAO, e.g. DESI Collaboration et al., 2024), and measurements of the luminosity distance ($D_{L}$) from supernovae (SNe, e.g. Riess et al., 2021). Other probes, such as redshift-space distortions (RSDs) (e.g. Scoccimarro, 2004), galaxy clustering (e.g. Cacciato et al., 2009), CMB lensing (e.g. Planck Collaboration et al., 2020b), and cosmic shear measure the properties of the large-scale structure (LSS) over a range of scales and with cosmic time. Cosmic shear is the distortion of galaxy shapes by weak gravitational lensing (WL) by the large-scale structure, and it is a powerful tool to map the matter distribution in the universe (for a review, see e.g. Kilbinger, 2015). Many of the LSS probes report values of the $S_{\mathrm{8}}$ parameter that are in mild tension with CMB predictions. This parameter is well constrained by WL inferences and measures a combination of the matter density and the amplitude of density fluctuations. The $S_{\mathrm{8}}$ tension manifests itself between CMB and cosmic shear (e.g. Hildebrandt et al., 2017; Asgari et al., 2021; Secco et al., 2022; Dark Energy Survey and Kilo-Degree Survey Collaboration et al., 2023), galaxy clustering (e.g. Tröster et al., 2020; Philcox & Ivanov, 2022), RSDs (e.g. Benisty, 2021), and thermal Sunyaev-Zel’dovich (tSZ) measurements (e.g. McCarthy et al., 2014, 2023).
@@
Clusters may be selected through different observables (e.g. optical richness, SZ or X-ray), but selection effects and model assumptions typically limit cluster cosmology inferences (e.g. Nagai et al., 2007; Kugel et al., 2024b). Whereas a Compton-y selected sample can provide cleaner selection effects, and more accurate mass estimates can be obtained by calibrating mass proxies using lensing measurements, the approach still suffers from not being able to definitively determine the completeness of the sample as clusters may be missing from the sample due to variations in their temperature and gas content (e.g. Kugel et al., 2024a). Additionally, current cluster inferences require several model assumptions, such as the choice of HMF, the assumed functional form for observable-mass scaling relations and the associated scatter, which may lead to biased results for next-generation measurements that aim to do precision cosmology at the sub-per cent level (Bocquet et al., 2016; Kugel et al., 2024b).
@@
Recently, Harnois-Deraps et al. (2024) carried out an inference using WL peaks selected on tomographic WL convergence maps from the KiDS-1000 and DES Y1 data releases. They found that this can greatly improve the results compared with selecting peaks from a single integrated WL convergence map as it can probe the evolution of the growth of structure, already highlighting the possibility of including redshift information. We aim to assign a redshift to every peak and thus expect to include more redshift information than in their study. Also, Chen et al. (2024) and Chiu et al. (2024) carried out a similar analysis as the idea explored in this paper by using HSC Y3 data and by selecting clusters through the peaks of WL aperture mass maps. Their sample only consists of 149 clusters selected through WL, and they thus cannot yet provide meaningful constraints. Nonetheless, it highlights the viability of the approach. For a recent review of peaks selected from aperture mass maps, see Oguri & Miyazaki (2024).
@@ 2 Theory
Lu & Haiman (2021) have argued that for a Stage III WL peak inference assuming the Born approximation leads to biased results, and a more sophisticated approach such as backward ray tracing should be applied. In this paper, we do not aim to carry out parameter inference but are instead interested in the origin and redshift evolution of a sample of high-valued WL peaks. Although the exact number of WL peaks at fixed WL convergence values may differ slightly depending on the map construction algorithm, we do not expect the origin of these peaks to change, as they should still stem from the same massive structures in the universe. As the Born approximation facilitates a more straightforward and computationally less expensive way to assign a redshift (as will become clear in Section 3.2), we adopt it in our analysis. Additionally, as we ignore several second-order effects such as lens-lens coupling (Bernardeau et al., 1997), and possible biasing factors such as dilution by cluster member contamination (Applegate et al., 2014; Medezinski et al., 2018) and source clustering (Gatti et al., 2023), the analysis is to some extent idealized.
@@ 3.3 Halo light-cone
To match the peaks to haloes, we use the FLAMINGO halo light-cones. FLAMINGO uses a modified version of the public halo finder Hierarchical Bound-Tracing (HBT+) to identify haloes and track their formation and evolution, including mergers (Han et al., 2018, Forouhar Moreno et al. in prep). The halo finder is used to identify central haloes and their most-bound BH particle, which is set as the centre of the halo. The 3-dimensional positions of the haloes are then converted to angular positions and redshifts, as seen by the different virtual observers in the runs, and saved as halo light-cone files. We directly use these files for our halo-peak matching. We rotate the angular coordinates of the haloes using the same rotations as the light-cone mass maps. On top of the halo finder, which determines the centres of the haloes in the simulation, the FLAMINGO data includes a large set of precomputed quantities in Spherical Overdensity APerture (SOAP) catalogues (McGibbon et al. in prep). From the SOAP catalogues, for each halo, we use $M_{\mathrm{200c}}$, the mass within a spherical aperture with radius $R_{\mathrm{200c}}$ enclosing an average overdensity of $\langle\rho(z)\rangle=200\rho_{\mathrm{crit}}(z)$, of the closest snapshot in redshift.
@@ 3.4 Halo - peak matching
To match the haloes to the peaks, we determine the great circle distance ($\Delta\theta$) between each peak and all nearby haloes. First, we project the haloes onto the HEALPix pixels according to their location on the halo light-cone. Then, using the healpy function query__\__disc, we select all haloes within 2 degrees of the peak. We then calculate the distance between the peak and all haloes in the angular aperture and scale the distances by the quantity $\theta_{\mathrm{200c}}=R_{\mathrm{200c}}/d_{\mathrm{A}}$, which is a proxy for the angular size of the halo on the sky, where $d_{\mathrm{A}}$ is the angular diameter distance. Then, we select the closest halo in $\Delta\theta/\theta_{\mathrm{200c}}$ for each peak. We do not require a bijective match between the halo and peak as we find that a single halo may cause multiple peaks in the final map, particularly for massive, low-redshift, or merging systems.
@@ 4.1 Peak distribution in halo mass
The bottom panel of Fig. 4 shows the fractional contribution of all mass bins. As expected, we find that, on average, higher-valued WL peaks correspond to more massive haloes. We find qualitatively similar results to Hamana et al. (2004); Yang et al. (2011); Liu & Haiman (2016), who found SNR > 3.5 peaks are predominantly caused by a single massive halo each. We can now study the peaks as a function of their halo mass in more detail. For $\kappa>0.25$, we find that over 95 per cent of the peaks can be attributed to haloes with $M_{\mathrm{200c}}>10^{14}$ $\mathrm{M_{\odot}}$. Of all SNR > 5 peaks, the majority are matched to haloes with $M_{\mathrm{200c}}=$$10^{14}-10^{14.5}$ $\mathrm{M_{\odot}}$. In Broxterman et al. (2024), we already hypothesised that the WL peaks originate primarily from $M_{\mathrm{200c}}>10^{14}$ $\mathrm{M_{\odot}}$ haloes based on the differences between the thermal and kinetic jet AGN variations and the comparison of the HMFs of these runs. We now confirm this hypothesis.
@@ 4.2.1 Comparison to X-ray and SZ cluster abundance inferences
To conclude, the SNR > 5 WL peak sample that Stage IV WL surveys will measure corresponds to massive clusters and has a purity and completeness equal to or better than state-of-the-art X-ray or SZ cluster abundance inferences. This especially holds for redshifts up to $z\approx 0.7$ and may be improved further by combining smoothing scales as illustrated in Appendix A. At the same time, WL peaks circumvent some of the selection effects and model assumptions, as both SZ and X-ray analyses require assumptions on the parametric form and scatter in the scaling relation relating the observable to the halo mass. Compared to an X-ray or SZ cluster analysis, the limiting factor of this analysis is that the WL kernel restricts the accessible redshift range. Therefore, the redshift range that we probe does not extend to as high a redshift as is probed by X-ray or SZ cluster analyses. In Section 4.1 we have shown that the SNR > 5 peaks probed by stage IV WL surveys primarily trace $M_{\mathrm{200c}}>10^{14}~{}\mathrm{M_{\odot}}$ haloes. Physically, these correspond to massive clusters of galaxies that should be visible in the optical, SZ, and X-ray (e.g. Allen et al., 2011). The analysis of this paper could, therefore, possibly be extended by including SZ-selected clusters for higher redshifts (e.g. Bocquet et al., 2023, 2024).
@@ 4.4 Baryonic and cosmology impact
Next, we compare the distributions of the virtual observers in the 1 cGpc simulations with different levels of baryonic feedback or cosmologies. We emphasize that all observers in the L1 boxes are located at the same position and that the simulations were initiated with the same initial phases. All SNR > 5 peaks in the final WL convergence maps are assigned the redshift of the lens plane that contributes most to the WL peak convergence value. At each lens plane, the WL peaks are matched to individual haloes and we assign the redshift of the haloes to the peaks. [^3] The FLAMINGO halo light-cones for L1__\__m9__\__DMO and the decaying dark matter variations are currently unavailable. We, therefore, show the results of these variations using the central redshifts of the shells to which the peaks have been assigned. As we only show relative differences to L1__\__m9, we also restrict the L1__\__m9 shell redshifts when comparing to L1__\__m9__\__DMO, PlanckDCDM12 and PlanckDCDM24. We confirmed that the ratio to L1__\__m9 is the same for the simulation variations that do have halo light-cones. We then adjust the redshifts to account for the expected uncertainty for the massive objects we probe (equation 11). In Fig. 9, we show the ratio of the redshift distributions ($\mathrm{d}n(z)/\mathrm{d}z$) of the WL peaks of all L1 variations with respect to L1__\__m9 (the green curve in Fig. 7) as a function of the perturbed observed redshifts $z_{\mathrm{obs}}$. However, we have found that the same conclusions hold when not including any observational redshift uncertainty or catastrophic outliers to the degree of the Euclid requirements. We bin the distributions using $\Delta z=0.1$.
@@
Figure 9: Ratio of the redshift distribution of the WL peak number density relative to the prediction for L1__\__m9 (green curve in Fig. 7). The left panel includes the variations varying the gas fractions in clusters, with more suppressed gas fractions indicated by an increasingly darker blue color and the DMO signal in black. The middle panel shows the remaining baryonic variations with a different AGN model (‘Jet’) or a suppressed galaxy stellar mass function (‘M*$-\sigma$’). The right panel shows the cosmology variations, including multiple variations using more massive neutrinos or decaying dark matter (‘DCDM’). The shaded green regions indicate the expected Poisson uncertainty on the Euclid DR1 and DR3 measurements. The baryonic variations (left and central panels) show little change in the shape of the redshift distribution compared to the fiducial model, while some of the cosmology variations, particularly the more massive neutrino variations and LS8, show a different shape redshift evolution. The cosmology variations also show a greater overall offset than the baryonic variations. The shape of the redshift distribution of high-SNR WL peaks is insensitive to the intensity of baryonic feedback, but is sensitive to cosmological parameters, making it potentially valuable for cosmological inference.
@@
The green shaded region indicates the Poisson error for the expected number of peaks in Euclid data release 1 (DR1, $\sim 2,500$ deg2) and for the final data release (DR3, $\sim 14,000$ deg2). The Poisson errors overestimate the differences between the simulation variations as all L1 variations were initiated with the same initial phases, and we thus expect the simulations to be correlated.
@@
The lack of differences in the redshift evolution between the baryonic feedback variations shows that the redshift distribution of the high-SNR WL peaks is set by the formation time of massive haloes and is insensitive to the strength of the baryonic feedback that manifests itself primarily at a later stage. As the shape of the L1__\__m9__\__DMO redshift evolution is consistent with that for the hydrodynamical simulations assuming the same cosmology, the redshift evolution of the WL peak sample studied in this work may be modeled with $N$-body simulations only.
Instead, if we focus on the cosmology variations in the right panel, we see greater differences in amplitude and evolution. We distinguish two subgroups in this panel. The first group, consisting of the neutrino mass variations and LS8, shows a different redshift evolution compared to the runs using the fiducial cosmology. Neutrinos act as a form of hot DM, thereby suppressing structure formation (e.g. Lesgourgues & Pastor, 2006). This is reflected in the lower value of $\sigma_{\mathrm{8}}$, as shown in the next to last column in Table 1. LS8, which has a lower value of $\sigma_{8}$ by design, and the higher neutrino mass variations have a lower value of $\sigma_{\mathrm{8}}$ than the fiducial cosmology. Consequently, there are fewer massive haloes in these simulations. This difference in the number of massive clusters in these simulations, compared to L1__\__m9, increases with redshift. This is also reflected in the redshift distributions, as those belonging to the more massive neutrino and LS8 variations are increasingly more suppressed at higher $z$.

### v2

@@ Abstract
Weak gravitational lensing (WL) convergence peaks contain valuable cosmological information in the regime of non-linear collapse. Using the FLAMINGO suite of cosmological hydrodynamical simulations, we study the physical origin and redshift distributions of the objects generating WL peaks selected from a WL convergence map mimicking a Euclid signal. We match peaks to individual haloes and show that the high signal-to-noise ratio (SNR > 5) WL peaks measured by Stage IV WL surveys primarily trace $M_{\mathrm{200c}}>10^{14}~{}\mathrm{M_{\odot}}$ haloes. We find that the WL peak sample can compete with the purity and completeness of state-of-the-art X-ray and Sunyaev-Zel’dovich cluster abundance inferences. By comparing the distributions predicted by simulation variations that have been calibrated to the observed gas fractions of local clusters and the present-day galaxy stellar mass function, or shifted versions of these, we illustrate that the shape of the redshift distribution of SNR > 5 peaks is insensitive to baryonic physics while it does change with cosmology. The difference highlights the potential of using WL peaks to constrain cosmology. As the WL convergence and redshift number densities of WL peaks scale differently with cosmology and baryonic feedback, WL peak statistics can simultaneously calibrate baryonic feedback and constrain cosmology.
@@ 1 Introduction
The cornerstone of modern cosmology is the $\Lambda$CDM model, consisting primarily of dark energy ($\Lambda$) and cold dark matter (CDM). The model can predict a wide range of cosmological observables, including the cosmic microwave background (CMB, e.g. Planck Collaboration et al., 2020a), baryonic acoustic oscillations (BAO, e.g. DESI Collaboration et al., 2025), and measurements of the luminosity distance ($D_{L}$) from supernovae (SNe, e.g. Riess et al., 2021). Other probes, such as redshift-space distortions (RSDs) (e.g. Scoccimarro, 2004), galaxy clustering (e.g. Cacciato et al., 2009), CMB lensing (e.g. Planck Collaboration et al., 2020b), and cosmic shear measure the properties of the large-scale structure (LSS) over a range of scales and with cosmic time. Cosmic shear is the distortion of galaxy shapes by weak gravitational lensing (WL) by the large-scale structure, and it is a powerful tool to map the matter distribution in the universe (for a review, see e.g. Kilbinger, 2015). Many of the LSS probes report values of the $S_{\mathrm{8}}$ parameter that are in mild tension with CMB predictions. This parameter is well constrained by WL inferences and measures a combination of the matter density and the amplitude of density fluctuations. The $S_{\mathrm{8}}$ tension manifests itself between CMB and cosmic shear (e.g. Hildebrandt et al., 2017; Asgari et al., 2021; Secco et al., 2022; Dark Energy Survey and Kilo-Degree Survey Collaboration et al., 2023), galaxy clustering (e.g. Tröster et al., 2020; Philcox & Ivanov, 2022), RSDs (e.g. Benisty, 2021), and thermal Sunyaev-Zel’dovich (tSZ) measurements (e.g. McCarthy et al., 2014, 2023).
@@
Clusters may be selected through different observables (e.g. optical richness, SZ or X-ray), but selection effects and model assumptions typically limit cluster cosmology inferences (e.g. Nagai et al., 2007; Kugel et al., 2025). Whereas a Compton-y selected sample can provide cleaner selection effects, and more accurate mass estimates can be obtained by calibrating mass proxies using lensing measurements, the approach still suffers from not being able to definitively determine the completeness of the sample as clusters may be missing from the sample due to variations in their temperature and gas content (e.g. Kugel et al., 2024). Additionally, current cluster inferences require several model assumptions, such as the choice of HMF, the assumed functional form for observable-mass scaling relations and the associated scatter, which may lead to biased results for next-generation measurements that aim to do precision cosmology at the sub-per cent level (Bocquet et al., 2016; Kugel et al., 2025).
@@
Recently, Harnois-Déraps et al. (2024) carried out an inference using WL peaks selected on tomographic WL convergence maps from the KiDS-1000 and DES Y1 data releases. They found that this can greatly improve the results compared with selecting peaks from a single integrated WL convergence map as it can probe the evolution of the growth of structure, already highlighting the possibility of including redshift information. We aim to assign a redshift to every peak and thus expect to include more redshift information than in their study. Also, Chen et al. (2025) and Chiu et al. (2024) carried out a similar analysis as the idea explored in this paper by using HSC Y3 data and by selecting clusters through the peaks of WL aperture mass maps. Their sample only consists of 149 clusters selected through WL, and they thus cannot yet provide meaningful constraints. Nonetheless, it highlights the viability of the approach. For a recent review of peaks selected from aperture mass maps, see Oguri & Miyazaki (2024).
@@ 2 Theory
Lu & Haiman (2021) have argued that for a Stage III WL peak inference assuming the Born approximation leads to biased results, and a more sophisticated approach such as backward ray tracing should be applied. In this paper, we do not aim to carry out parameter inference but are instead interested in the origin and redshift evolution of a sample of high-valued WL peaks. Although the exact number of WL peaks at fixed WL convergence values may differ slightly depending on the map construction algorithm, we do not expect the origin of these peaks to change, as they should still stem from the same massive structures in the universe. As the Born approximation facilitates a more straightforward and computationally less expensive way to assign a redshift (as will become clear in Section 3.2), we adopt it in our analysis. Additionally, as we ignore several second-order effects such as lens-lens coupling (Bernardeau et al., 1997), and possible biasing factors such as dilution by cluster member contamination (Applegate et al., 2014; Medezinski et al., 2018), the analysis is to some extent idealized. Some of these effects, such as source clustering, are expected to have a minor impact on WL peak statistics (Gatti et al., 2023). For others, such as galaxy intrinsic alignment, the impact, although small, is non-negligible (Zhang et al., 2022). As the impact of these effects is less than 1 per cent on the power spectrum (Kilbinger, 2015), we do not expect them to impact the conclusions of this work as we study the relative differences between the different simulation variations, but an analysis using legacy Stage IV WL survey data should consider them in more detail.
@@ 3.3 Halo light-cone
To match the peaks to haloes, we use the FLAMINGO halo light-cones. FLAMINGO uses the halo finder HBT-HERONS to identify haloes and track their formation and evolution, including mergers (Han et al., 2018; Forouhar Moreno et al., 2025). The halo finder is used to identify central haloes and their most-bound BH particle, which is set as the centre of the halo. The 3-dimensional positions of the haloes are then converted to angular positions and redshifts, as seen by the different virtual observers in the runs, and saved as halo light-cone files. We directly use these files for our halo-peak matching. We rotate the angular coordinates of the haloes using the same rotations as the light-cone mass maps. On top of the halo finder, which determines the centres of the haloes in the simulation, the FLAMINGO data includes a large set of precomputed quantities in Spherical Overdensity APerture (SOAP) catalogues (McGibbon et al. in prep). From the SOAP catalogues, for each halo, we use $M_{\mathrm{200c}}$, the mass within a spherical aperture with radius $R_{\mathrm{200c}}$ enclosing an average overdensity of $\langle\rho(z)\rangle=200\rho_{\mathrm{crit}}(z)$, of the closest snapshot in redshift.
@@ 3.4 Halo - peak matching
To match the haloes to the peaks, we determine the great circle distance ($\Delta\theta$) between each peak and all nearby haloes. First, we project the haloes onto the HEALPix pixels according to their location on the halo light-cone. Then, using the healpy function query__\__disc, we select all haloes within 2 degrees of the peak. We then calculate the distance between the peak and all haloes in the angular aperture and scale the distances by the quantity $\theta_{\mathrm{200c}}=R_{\mathrm{200c}}/d_{\mathrm{A}}$, which is a proxy for the angular size of the halo on the sky, where $d_{\mathrm{A}}$ is the angular diameter distance. Then, we select the closest halo in $\Delta\theta/\theta_{\mathrm{200c}}$ for each peak. [^3] We find the results do not change if we instead match peaks to haloes based on $\Delta\theta/L_{*,50\mathrm{kpc}}$ or $\Delta\theta/M_{*,50\mathrm{kpc}}$, where $L_{*,50\mathrm{kpc}}$ and $M_{*,50\mathrm{kpc}}$ are the stellar luminosity in the GAMA $r$ band and stellar mass in a 50 kpc aperture, respectively. We do not require a bijective match between the halo and peak as we find that a single halo may cause multiple peaks in the final map, particularly for massive, low-redshift, or merging systems. If, instead, we match peaks to the closest halo in $\Delta\theta/\theta_{\mathrm{200c}}$ without first assigning the peak to a redshift, i.e. matching to all haloes in the same angular region we find that 95 (97) per cent of the SNR $>$ 5 (8) peaks are matched to the same haloes. The comparison indicates minor dependence on the matching algorithm due to projection effects but we find this does not impact our main conclusions.
@@ 4.1 Peak distribution in halo mass
The bottom panel of Fig. 4 shows the fractional contribution of all mass bins. As expected, we find that, on average, higher-valued WL peaks correspond to more massive haloes. We find qualitatively similar results to Hamana et al. (2004); Yang et al. (2011); Liu & Haiman (2016), who found SNR > 3.5 peaks are predominantly caused by a single massive halo each. We can now study the peaks as a function of their halo mass in more detail. For $\kappa>0.25$, we find that over 95 per cent of the peaks can be attributed to haloes with $M_{\mathrm{200c}}>10^{14}$ $\mathrm{M_{\odot}}$. Of all SNR > 5 peaks, the majority are matched to haloes with $M_{\mathrm{200c}}~{}=~{}10^{14}-10^{14.5}$ $\mathrm{M_{\odot}}$. In Broxterman et al. (2024), we already hypothesised that the WL peaks originate primarily from $M_{\mathrm{200c}}>10^{14}$ $\mathrm{M_{\odot}}$ haloes based on the differences between the thermal and kinetic jet AGN variations and the comparison of the HMFs of these runs. We now confirm this hypothesis.
@@ 4.2.1 Comparison to X-ray and SZ cluster abundance inferences
To conclude, the SNR > 5 WL peak sample that Stage IV WL surveys will measure corresponds to massive clusters and has a purity and completeness equal to or better than state-of-the-art X-ray or SZ cluster abundance inferences. This especially holds for redshifts up to $z\approx 0.7$ and may be improved further by combining smoothing scales as illustrated in Appendix A. At the same time, WL peaks circumvent some of the selection effects and model assumptions, as both SZ and X-ray analyses require assumptions on the parametric form and scatter in the scaling relation relating the observable to the halo mass. Compared to an X-ray or SZ cluster analysis, the limiting factor of this analysis is that the WL kernel restricts the accessible redshift range. Therefore, the redshift range that we probe does not extend to as high a redshift as is probed by X-ray or SZ cluster analyses. In Section 4.1 we have shown that the SNR > 5 peaks probed by stage IV WL surveys primarily trace $M_{\mathrm{200c}}>10^{14}~{}\mathrm{M_{\odot}}$ haloes. Physically, these correspond to massive clusters of galaxies that should be visible in the optical, SZ, and X-ray (e.g. Allen et al., 2011). The analysis of this paper could, therefore, possibly be extended by including SZ-selected clusters for higher redshifts (e.g. Bocquet et al., 2024a, b).
@@ 4.4 Baryonic and cosmology impact
Next, we compare the distributions of the virtual observers in the 1 cGpc simulations with different levels of baryonic feedback or cosmologies. We emphasize that all observers in the L1 boxes are located at the same position and that the simulations were initiated with the same initial phases. All SNR > 5 peaks in the final WL convergence maps are assigned the redshift of the lens plane that contributes most to the WL peak convergence value. At each lens plane, the WL peaks are matched to individual haloes and we assign the redshift of the haloes to the peaks. [^4] The FLAMINGO halo light-cones for L1__\__m9__\__DMO and the decaying dark matter variations are currently unavailable. We, therefore, show the results of these variations using the central redshifts of the shells to which the peaks have been assigned. As we only show relative differences to L1__\__m9, we also restrict the L1__\__m9 shell redshifts when comparing to L1__\__m9__\__DMO, PlanckDCDM12 and PlanckDCDM24. We confirmed that the ratio to L1__\__m9 is the same for the simulation variations that do have halo light-cones. We then adjust the redshifts to account for the expected uncertainty for the massive objects we probe (equation 11). In Fig. 9, we show the ratio of the redshift distributions ($\mathrm{d}n(z)/\mathrm{d}z$) of the WL peaks of all L1 variations (top panels) and their ratio with respect to L1__\__m9 (bottom panels) as a function of the perturbed observed redshifts $z_{\mathrm{obs}}$. However, we have found that the same conclusions hold when not including any observational redshift uncertainty or catastrophic outliers to the degree of the Euclid requirements. We bin the distributions using $\Delta z=0.1$.
@@
Figure 9: Top: number density of the redshifts of SNR > 5 WL peaks for all L1__\__m9 FLAMINGO variations. Bottom: Ratio of the redshift distribution of the WL peak number density relative to the prediction for L1__\__m9 (dark green curve). The left panels includes the variations varying the gas fractions in clusters, with more suppressed gas fractions indicated by an increasingly darker blue colour and the DMO signal in black. The middle panels shows the remaining baryonic variations with a different AGN model (‘Jet’) or a suppressed galaxy stellar mass function (‘M*$-\sigma$’). The right panels shows the cosmology variations, including multiple variations using more massive neutrinos or decaying dark matter (‘DCDM’). The shaded green regions in the bottom panels indicate the expected Poisson uncertainty on the Euclid DR1 and DR3 measurements. The baryonic variations (left and central panels) show little change in the shape of the redshift distribution compared to the fiducial model, while some of the cosmology variations, particularly the more massive neutrino variations and LS8, show a different shape redshift evolution. The cosmology variations also show a greater overall offset than the baryonic variations. The shape of the redshift distribution of high-SNR WL peaks is insensitive to the intensity of baryonic feedback, but is sensitive to cosmological parameters, making it potentially valuable for cosmological inference.
@@
The green shaded region in the bottom panels indicates the Poisson error for the expected number of peaks in Euclid data release 1 (DR1, $\sim 2,500$ deg2) and for the final data release (DR3, $\sim 14,000$ deg2). The Poisson errors overestimate the differences between the simulation variations as all L1 variations were initiated with the same initial phases, and we thus expect the simulations to be correlated.
@@
The lack of differences in the redshift evolution between the baryonic feedback variations is consistent with the interpretation that the redshift distribution of the high-SNR WL peaks is set by the formation time of massive haloes and is insensitive to the strength of the baryonic feedback that manifests itself primarily only once a halo has formed and has accreted sufficient mass. The strength of the feedback does not impact the time at which a halo has formed but only whether it creates a WL signal significant enough to be included in the SNR > 5 sample, as reflected in the amplitude differences as shown in Figure 9. This reasoning still holds if baryonic feedback is redshift dependent or if haloes of the same mass experience different baryonic feedback at different redshifts, provided the WL peaks correspond to single massive objects, as we have shown is the case, and the baryonic evolution is independent of cosmology, which is to first order the case for WL peaks measured in the limited number of FLAMINGO variations at different cosmologies. As the shape of the L1__\__m9__\__DMO redshift evolution is consistent with that for the hydrodynamical simulations assuming the same cosmology, the redshift evolution of the WL peak sample studied in this work may be modeled with $N$-body simulations only.
Instead, if we focus on the cosmology variations in the right panels, we see greater differences in amplitude and evolution. We distinguish two subgroups in these panels. The first group, consisting of the neutrino mass variations and LS8, shows a different redshift evolution compared to the runs using the fiducial cosmology. Neutrinos act as a form of hot DM, thereby suppressing structure formation (e.g. Lesgourgues & Pastor, 2006). This is reflected in the lower value of $\sigma_{\mathrm{8}}$, as shown in the next to last column in Table 1. LS8, which has a lower value of $\sigma_{8}$ by design, and the higher neutrino mass variations have a lower value of $\sigma_{\mathrm{8}}$ than the fiducial cosmology. Consequently, there are fewer massive haloes in these simulations. This difference in the number of massive clusters in these simulations, compared to L1__\__m9, increases with redshift. This is also reflected in the redshift distributions, as those belonging to the more massive neutrino and LS8 variations are increasingly more suppressed at higher $z$.